OpenStudy (anonymous):

Help!!! If two electrical resistors with resistance R1 and R2 are connected, then the total resistance in the circuit is given by the complex rational expression (see reply for the complex rational express) Simplify the exression Now find the total resistance if R1=10 ohms and R2=20 ohms

OpenStudy (anonymous):

\[\frac{ 1 }{ \frac{ 1 }{ R1 }+ \frac{ 1 }{ R2 } }\]

OpenStudy (anonymous):

@whpalmer4 do you think you can help?

OpenStudy (whpalmer4):

It's just straightforward algebra, and you know the values, what's the issue?

OpenStudy (cggurumanjunath):



OpenStudy (cggurumanjunath):